Assume that we want to construct a confidence interval. Do one of the​ following, as​ appropriate: (a) find the critical value

t Subscript alpha divided by 2tα/2​,

​(b) find the critical value z Subscript alpha divided by 2zα/2​,
or​

(c) state that neither the normal distribution nor the t distribution applies.The confidence level is

90​%,

sigmaσ is not​ known, and the histogram of 57 player salaries​ (in thousands of​ dollars) of football players on a team is as shown.
Answer #1

Solution :

Given that,

= 1 - 0.90 = 0.10

/ 2 = 0.05

n = 57

df = 57 - 1 = 56

t 0.05,54 = 1.674